OpenStudy (xmoses1):

If p∨q is true, then what must be true about the truth values of p and q? both are true both are false at least one is true

OpenStudy (raden):

at least one is true

OpenStudy (xmoses1):

thank you! i have one more

OpenStudy (xmoses1):

Two angles are complementary. One of the angles is twice the other. The larger angle has a measure of 120° 60° 45°

OpenStudy (xmoses1):

I am not sure how to do this problem. I know that complimentary means that the angles add up to 90 degrees.

OpenStudy (raden):

let one angle is x and the other 2x x+2x = 90 solve for x

OpenStudy (xmoses1):

Thirty degrees?

OpenStudy (xmoses1):

Oh it is 45!

OpenStudy (raden):

yes, x = 30 obviously as the the larger is 2 * 30 = 60

OpenStudy (xmoses1):

So it isnt 45 degrees?

OpenStudy (raden):

no, it is 60 (twice of x)

OpenStudy (raden):

with x is the first angle

OpenStudy (xmoses1):

ok, the reason that i thought that is because : x+2x = 90 I divided both 2x and 90 by two and was left over with x=45. But i see that you are right:)

OpenStudy (xmoses1):

Thank You!

OpenStudy (raden):

x+2x=3x so, we have 3x = 90 divide by 3 to both sides x = 90/3 = 30 you're welcome

OpenStudy (xmoses1):

Thank you for clarifying how to do this problem:)